OpenStudy (anonymous):

-a^3 - 100a would this be considered prime? meaning I could not factor it

OpenStudy (anonymous):

beat me to it

OpenStudy (anonymous):

\[-a^3 - 100a = -a(a^2 + 100)\]THis is far as you can actually factor it.

OpenStudy (anonymous):

oh ok.. thank you

OpenStudy (anonymous):

so I wouldnt have a common factor?

OpenStudy (anonymous):

or would it be a

OpenStudy (anonymous):

Well, you can factor out a GCF which would be -a.

OpenStudy (anonymous):

so when worked out it looks like this: -a^3 - 100a -a(a^2+100) with a GCF of -a

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\[-(a)(a+10i)(a-10i)\]The *complete* factorization.

OpenStudy (anonymous):

Yes. True. I just didn't do that because I'm not sure if you learned i yet.

OpenStudy (anonymous):

this would be considered a perfect square

OpenStudy (anonymous):

wouldnt the (a-10) be the same on both sides due to the -100a

OpenStudy (anonymous):

What? Could you demonstrate what you mean?

OpenStudy (anonymous):

(-a)(a+10)(a+10) because the positive and the negative will equal the -100a

OpenStudy (anonymous):

nevermind they would both be (-)

OpenStudy (anonymous):

Let's try to multiply all of this together then and see what we get? \[(-a)(a^2 + 20a + 100) \rightarrow -a^3 - 20a^2 - 100a\]

OpenStudy (anonymous):

See?

OpenStudy (anonymous):

I mean (+)

OpenStudy (anonymous):

What I wrote was correct. Don't worry about it.

OpenStudy (anonymous):

-a^3-100a -a(a^2 +100) (-a)(a+10)(a+10) when multiplied back (-a) (a^2 + 100) a^2 - 100a Right?

OpenStudy (anonymous):

What? Not quite...

OpenStudy (anonymous):

a^3 - 100a

OpenStudy (anonymous):

\[(a + 10)(a + 10) = a(a + 10) + 10(a + 10) = a^2 + 20a + 100\]Not \(a^2 + 100\)

OpenStudy (anonymous):

where did the 20a come from?

OpenStudy (anonymous):

You have to FOIL it. \[a \times a = a^2\]\[a \times 10 = 10a\]\[10 \times a = 10a\]\[10 \times 10 = 100\]

OpenStudy (anonymous):

so how do I come up with a^3

OpenStudy (anonymous):

when multiplied by the 10a?

OpenStudy (anonymous):

sorry I am completely lost.. I thought I had it, but apparently I dont :(

OpenStudy (anonymous):

You end up with something that isn't the correct answer which is what you were saying is correct.

OpenStudy (anonymous):

Look. You sai that -a(a + 10)(a + 10) = -a^3 - 100a right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If you look at what I said above, (a + 10)(a + 10) = a^2 + 20a + 100

OpenStudy (anonymous):

ok but I thought you had to multiple the (a +10)(a +10)

OpenStudy (anonymous):

The: \[-a(a^2 + 20a + 100) = -a^2 - 20a^2 - 100a\]See?

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

Are you confused about the (a + 10)(a + 10) part?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Ok. Look. When you multiply two binomials together, you do something called FOIL. First Outside Inside Last Multiply the terms in that order. First terms are a and a. (a)(a) = a^2 Outside Terms: (a)(10) = 10a Inside Terms: (10)(a) = 10a Last Terms: (10)(10) = 100 Then you add them together. a^2 + 10a + 10a + 100 = a^2 + 20a + 100

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

ok so when I multiply back to check the answer I am not getting the original binomial. Am I not supposed to?

OpenStudy (anonymous):

You're supposed to get the original answer back. THat's why I said that after \[-a(a^2 + 100)\]is unfactorable because you don't get that back.

OpenStudy (anonymous):

So \[-a(a^2 + 100)\]is as far as you can get.

OpenStudy (anonymous):

ok. so I have confused myself once again. making this harder than it really needs to be sorry.

OpenStudy (anonymous):

Don't worry :) You're fine

OpenStudy (anonymous):

this binomial would not be considered square correct?

OpenStudy (anonymous):

Yes. A perfect square trinomial is in the form: \[(a + b)^2 = a^2 + 2ab + b^2\]\[(a - b)^2 = a^2 - 2ab + b^2\]

OpenStudy (anonymous):

ok thank you. was just checking

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

-a^3 – 100a The original expression to be factored. We notice that the leading coefficient is (–a) and the constant term is an even number. -a(a^2 – 100) The binomial cannot be factored any further and is considered complete. Would this explanation be correct?

OpenStudy (anonymous):

am I missing a step in the middle?

OpenStudy (anonymous):

Well, I would say that the justification for the first step is wrong. The original expression has a Greatest Common Factor that can be factored out. Noticing how -a is in bother terms, I factored it out.

OpenStudy (anonymous):

ok. the GCF would be the -a?

OpenStudy (anonymous):

or would it actually be 10